### v2

# The number of polygons inscribed into one and circumscribed around another polygon

Abstract
In the paper Poncelet type problem about the number of convex $n$-gons inscribed into one convex $n$-gon and circumscribed around another convex $n$-gon, is studied. It is proved that this number is at most 4. This contrasts with Poncelet type porisms where usually infinitude of such polygons is proved, provided that one polygon exists. An inequality involving ratio of lengths of line segments is used. Alternative way of using Maclaurin-Braikenridge’s conic generation method is also discussed. Properties related to constructibility with straightedge and compass are also studied. A new proof, based on mathematical induction, of generalized Maclaurin-Braikenridge’s theorem is given.
2020 Mathematics Subject Classification: Primary 51M04; Secondary 51M15, 51M16, 52C05.Key words and phrases: porism, Poncelet type theorems, inscribed polygon, circumscribed polygon, Maclaurin’s conic generation, Braikenridge’s theorem, geometric inequality.
1 Introduction
Poncelet in his introduction to famous porism theorem, studied momentarily the variable polygons having all but one of their vertices on sides of a given polygon and all the sides passing through the vertices of another polygon. He shows that the free vertex describes a conic ([21], Tome I, p. 308-314). Poncelet cites Brianchon [6], who cites Maclaurin [16] and Braikenridge [5] for the case of triangles $(n=3)$ (see [17] for the history). Poncelet started to work on these questions in 1813 when he was a prisoner of war in Russia. In Poncelet porisms infinitude of polygons inscribed into one ellipse and and circumscribed around another ellipse, given that one such polygon exists, is proved [21], [22] (see also [7], [8], [14]). This result inspired discovery of many other porisms commonly known as Poncelet type theorems. In Steiner porism infinitude of chains of tangent circles all of which are internally tangent to one circle and externally tangent to another circle, again provided that one such chain exists, is studied (see[20], p.98; [9], sect. 5.8). In Emch’s porism Steiner’s porism is generalized to chain of intersecting circles, where these intersections are on another circle [13] (see also [2]). In Zig-zag type porisms infinitude of equilateral polygons with vertices taken alternately on two circles (or lines), is studied [4], [15], [11]. Money-Coutts theorem (see [25] and its references) is about chain of tangent circles alternately inscribed into angles of a triangle. Generalizations and connections of these results with each other, were studied in [3], [23], [12]. Surprisingly, the original configuration of Maclaurin about polygons inscribed into one polygon and circumscribed around another polygon did not attract much attention after these powerful generalizations. In the current paper we will study in detail the case of convex polygons inscribed into one convex polygon and circumscribed around another one. The results are surprising as we will show that the maximal number of such polygons is not dependent on the number of sides of the polygons. Possibility of construction of such regular polygons with Euclidean instruments (straightedge and compass) is disscussed in Appendix A. At the end of the paper the case of generalized polygons, considered as collections of vertices or lines is also considered. For completeness we also provided proof of generalized Maclaurin-Braikenridge’s conic generation method in Appendix B.
2 Main results
Suppose polygon $A_{0}A_{1}A_{2}\ldots A_{n-1}$ $(n\geq 3$) is given. Choose a point on each side of this polygon. Denote these points by $B_{i}$, where $B_{i}\in A_{i}A_{i+1}$, $B_{i}\neq A_{i},A_{i+1}$ $(i=0,1,\ldots,n-1)$. For such polygons we say $B_{0}B_{1}B_{2}\ldots B_{n-1}$ is inscribed into $A_{0}A_{1}A_{2}\ldots A_{n-1}$ and $A_{0}A_{1}A_{2}\ldots A_{n-1}$ is circumscribed around $B_{0}B_{1}B_{2}\ldots B_{n-1}$. In the same way choose on each side of polygon $B_{0}B_{1}B_{2}\ldots B_{n-1}$ a point $C_{i}\in B_{i}B_{i+1}$ $(i=0,1,\ldots,n-1)$ and draw polygon $C_{0}C_{1}C_{2}\dots C_{n-1}$. Now fix polygons $A_{0}A_{1}A_{2}\ldots A_{n-1}$ and $C_{0}C_{1}C_{2}\dots C_{n-1}$. We want to find the maximum of the number of polygons like $B_{0}B_{1}B_{2}\ldots B_{n-1}$, which are inscribed into $A_{0}A_{1}A_{2}\ldots A_{n-1}$ and circumscribed around $C_{0}C_{1}C_{2}\dots C_{n-1}$. We will prove that this number is 4. Our strategy is to give an example where number 4 is realized and then prove that this number can not exceed 4.
Figure 1: How many convex polygons are there, which are inscribed into one convex polygon and circumscribed around another convex polygon?
The key ingredients of the proof will be the following example and lemma, which are interesting on their own.
Example 2.1.
Let $A_{0}A_{1}A_{2}\ldots A_{n-1}$ be a regular polygon with with side length equal to $2a$ and center $O$. Let $E_{i}$ be the midpoints of $A_{i}A_{i+1}$ $(i=0,1,\ldots,n-1)$. Let $B_{i}\in A_{i}E_{i}$ and denote $|A_{i}B_{i}|=x$ $(i=0,1,\ldots,n-1)$. Then $|E_{i}B_{i}|=a-x$ (see Fig. 2). Denote intersection point of line segments $OE_{i}$ and $B_{i}B_{i+1}$ by $C_{i}$. Then
$|C_{i}F_{i}|=\frac{x(a-x)\sin{\frac{2\pi}{n}}}{2a-2x\sin^{2}{\frac{\pi}{n}}}.$
Figure 2: Example 2.1.
Denote $f(x)=\frac{x(a-x)\sin{\frac{2\pi}{n}}}{2a-2x\sin^{2}{\frac{\pi}{n}}}.$ One can check that
$f\left(\frac{a}{2}\right)=f\left(\frac{a}{1+\cos^{2}{\frac{\pi}{n}}}\right)=\frac{a\sin{\frac{2\pi}{n}}}{4\left(1+\cos^{2}{\frac{\pi}{n}}\right)}.$
Therefore the two regular polygons $B_{0}B_{1}B_{2}\ldots B_{n-1}$ corresponding to $x=\frac{a}{2}$, $x=\frac{a}{1+\cos^{2}{\frac{\pi}{n}}}$ and their reflections with respect to line $OE_{i}$ form 4 polygons, which are inscribed into $A_{0}A_{1}A_{2}\ldots A_{n-1}$ and circumscribed around $C_{0}C_{1}C_{2}\dots C_{n-1}$ (see Fig. 3 for case $n=3$). Note that points $B_{i}$ corresponding to $x=\frac{a}{2}$ and $x=\frac{a}{1+\cos^{2}{\frac{\pi}{n}}}$ are constructible with compass and straightedge, provided that regular polygon $A_{0}A_{1}A_{2}\ldots A_{n-1}$ is given (see Appendix A).
Figure 3: Case $n=3$ for Example 2.1. Four triangles inscribed into one equilateral triangle and circubscribed around another one.
Lemma 2.2.
Let cevians $BD$ and $CE$ of $\triangle ABC$ intersect at point $G$. A line through point $G$ inersects line segments $BE$ and $CD$ at points $F$ and $H$. Then $\frac{|BF|}{|FE|}>\frac{|DH|}{|HC|}$.
Figure 4: Lemma 2.2: $\frac{|BF|}{|FE|}>\frac{|DH|}{|HC|}$.
Proof.
Let the line through point $E$ and parallel to line $AC$ intersect line segments $FG$ and $BG$ at points $I$ and $J$. By Menelaus’ theorem
$\frac{|BF|}{|FE|}\cdot\frac{|EI|}{|IJ|}\cdot\frac{|JG|}{|GB|}=1.$
By similarity of $\triangle EIG$ and $\triangle CHG$, $\triangle JIG$ and $\triangle DHG$, $\triangle EGJ$ and $\triangle CGD$, we obtain $\frac{|EI|}{|IJ|}=\frac{|CH|}{|HD|}$. Finally, noting $|JG|<|GB|$, we obtain
$\frac{|BF|}{|FE|}=\frac{|IJ|}{|EI|}\cdot\frac{|GB|}{|JG|}=\frac{|DH|}{|HC|}\cdot\frac{|GB|}{|JG|}>\frac{|DH|}{|HC|}.$
Theorem 2.3.
The maximum of the number of convex polygons $B_{0}B_{1}B_{2}\ldots B_{n-1}$, that can be inscribed into a given convex polygon $A_{0}A_{1}A_{2}\ldots A_{n-1}$ and circumscribed around around another given convex polygon $C_{0}C_{1}C_{2}\dots C_{n-1}$ is 4.
Proof.
We saw in Example 1.1 that the case with 4 polygons is possible. Suppose on the contrary that there is a configuration with at least 5 polygons. Denote 5 of them by $B^{j}_{0}B^{j}_{1}B^{j}_{2}\ldots B^{j}_{n-1}$ $(j=1,2,\ldots,5)$ so that $|A_{0}B^{j}_{0}|<|A_{0}B^{j+1}_{0}|$, for $j=1,2,3,4$. Immediately, $|A_{i}B^{j}_{i}|<|A_{i}B^{j+1}_{i}|$, for all $i=0,1,\ldots,n-1$ and $j=1,2,3,4$. Consider the first two of the polygons: $B^{1}_{0}B^{1}_{1}B^{1}_{2}\ldots B^{1}_{n-1}$ and $B^{2}_{0}B^{2}_{1}B^{3}_{2}\ldots B^{3}_{n-1}$. Then the vertices of $C_{0}C_{1}C_{2}\dots C_{n-1}$ are $C_{i}=B^{1}_{i}B^{1}_{i+1}\cap B^{2}_{i}B^{2}_{i+1}$ $(i=0,1,\ldots,n-1)$, where all the indices are taken modulo $n$. Note that $B^{3}_{0}B^{3}_{1}$ can not pass through $C_{0}$. Indeed, otherwise $B^{3}_{i}B^{3}_{i+1}$ pass through $C_{i}$ for all $i=0,1,\ldots,n-1$, and by Lemma 2.2,
$\frac{|B^{1}_{i}B^{2}_{i}|}{|B^{2}_{i}B^{3}_{i}|}>\frac{|B^{1}_{i+1}B^{2}_{i+1}|}{|B^{2}_{i+1}B^{3}_{i+1}|},$
for all $i=0,1,\ldots,n-1$, which is impossible, because it implies $\frac{|B^{1}_{0}B^{2}_{0}|}{|B^{2}_{0}B^{3}_{0}|}>\frac{|B^{1}_{0}B^{2}_{0}|}{|B^{2}_{0}B^{3}_{0}|}.$ So, by convexity of $A_{0}A_{1}A_{2}\ldots A_{n-1}$ and $C_{0}C_{1}C_{2}\dots C_{n-1}$, $B^{3}_{0}B^{3}_{1}$ passes through $C_{1}$. Similarly, $B^{4}_{0}B^{4}_{1}$, $B^{5}_{0}B^{5}_{1}$ pass through $C_{1}$. Then $B^{3}_{i}B^{3}_{i+1}$, $B^{4}_{i}B^{4}_{i+1}$, and $B^{5}_{i}B^{5}_{i+1}$ pass through $C_{i+1}$ for all $i=0,1,\ldots,n-1$. By Lemma 2.2,
$\frac{|B^{3}_{i}B^{4}_{i}|}{|B^{4}_{i}B^{5}_{i}|}>\frac{|B^{3}_{i+1}B^{4}_{i+1}|}{|B^{4}_{i+1}B^{5}_{i+1}|},$
for all $i=0,1,\ldots,n-1$, which is again impossible. This shows that our assumption was wrong and therefore the maximal number of polygons is 4. ∎
Figure 5: Proof of Theorem 2.3.
If $n>3$, then all sets
$\mathbb{M}_{i}=\triangle A_{i}A_{i+1}A_{i+2}\cap\triangle A_{i-1}A_{i}A_{i+1}\ (i=0,1,\ldots,n-1)$
are disjoint (only the interior regions of the triangles are considered). When the case of 4 inscribed and circumscribed polygons is realized, $C_{i}\in\mathbb{M}_{i}$. If $n=3$, then $C_{i}\in\mathbb{M}_{i}$ is trivially true, because $\triangle A_{i}A_{i+1}A_{i+2}$ $(i=0,1,2)$ coincide with $\triangle A_{0}A_{1}A_{2}$ and therefore $\mathbb{M}_{i}=\triangle A_{0}A_{1}A_{2}$.
3 Maclaurin-Braikenridge’s method of conic generation
It is well known that if triangles $A_{0}A_{1}A_{2}$ and $C_{0}C_{1}C_{2}$ are given, moving line $B_{0}B_{2}$ passes through $C_{2}$, where $B_{0}$ and $B_{2}$ are on lines $A_{0}A_{1}$ and $A_{0}A_{2}$, respectively, then the locus of intersection point $X$ of $B_{0}C_{0}$ and $B_{2}C_{1}$ is in general a conic (see Fig. 6). See e.g. [24], p. 230, [18], p. 22, [19], p. 97, [28], [20], p. 328, where this result, known as Maclaurin-Braikenridge’s method of conic generation, is generalized for polygons $A_{0}A_{1}A_{2}\ldots A_{n-1}$ and $C_{0}C_{1}C_{2}\dots C_{n-1}$. If $n=3$, then this conic passes through $A_{0},\ C_{0},\ C_{1},P,$ and $Q$, where $P=A_{0}A_{1}\cap C_{1}C_{2}$ and $Q=A_{0}A_{2}\cap C_{0}C_{2}$. This conic can intersect $A_{1}A_{2}$ at maximum 2 points shown in Figure 6 as $B^{1}_{1}$ and $B^{2}_{1}$, which can be starting point for the construction of 2 triangles $B^{1}_{0}B^{1}_{1}B^{1}_{3}$ and $B^{2}_{0}B^{2}_{1}B^{2}_{3}$ inscribed into $\triangle A_{0}A_{1}A_{2}$ and circumscribed around $\triangle C_{0}C_{1}C_{2}$. Therefore the key argument of using Lemma 2.2 in the proof of Theorem 2.3, can also be replaced by the fact that a general conic and a line can intersect at maximum 2 points.
Figure 6: Maclaurin-Braikenridge’s method of conic generation.
The generalization of Maclaurin-Braikenridge’s method was studied through analytic tools in [21], Tome I, p. 308-314 (see [6] for geometric consideration). The proof of case $n=3$ is the converse of Pascal’s theorem [9], which states that hexagon $A_{0}PC_{1}XC_{0}Q$ is inscribed into a conic ([26], p.321, [27], p. 377). The general case can be proved by induction on $n$ (see Appendix B). Using this general result it is possible to show that if the condition of convexity is lifted and the polygons are extended to include also the lines containing the sides of the polygons, then in the general case, the maximal number of polygons inscribed into one polygon and circumscribed around another one, can be as large as $n!(n-1)!$ (excuding degenerate cases when the conics can coincide with the lines giving infinitely many such polygons [21], Tome 2, p. 10, [19], p. 97). Indeed, there are $(n-1)!$ cyclic permutations of the sides of polygon $A_{0}A_{1}A_{2}\ldots A_{n-1}$. Between these sides the vertices of polygon $C_{0}C_{1}C_{2}\dots C_{n-1}$ can be put in $n!$ ways. Because of independence, we multiply them to obtain $n!(n-1)!$, then divide by 2 (orientation is not important) and finally multiply by 2 (each conic gives at most two polygons) to obtain again $n!(n-1)!$. For example, if $n=3$, then there are at most 12 triangles, all of which are shown in Fig. 7.
Figure 7: There are $3!2!=12$ inscribed/circumscribed triangles if $n=3$.
4 Appendix A
Suppose that regular polygon $A_{0}A_{1}A_{2}\ldots A_{n-1}$ is given. Note that point $B_{i}$ corresponding to $x=\frac{a}{2}$ is the midpoint of $E_{i}A_{i}$, which is easy to construct using unmarked ruler and compass. Below we will show how to construct using the same instruments point $B_{i}$ corresponding to $x=\frac{a}{1+\cos^{2}{\frac{\pi}{n}}}$. Let $E_{i}$ be, as before, the midpoint of side $A_{i}A_{i+1}$ of regular polygon $A_{0}A_{1}A_{2}\ldots A_{n-1}$.
  1. 1.
    Drop perpendicular $E_{i}C$ to line $OA_{i}$.
  2. 2.
    Drop perpendicular $CD$ to line $A_{i}A_{i+1}$.
  3. 3.
    Draw circle with center $A_{i+1}$ through $E_{i}$. Let this circle intersect line $A_{i+2}A_{i+1}$ outside of line segment $A_{i+2}A_{i+1}$ at point $G$.
  4. 4.
    Draw line through $E_{i}$ parallel to $DG$. Let this line intersect line $A_{i+2}A_{i+1}$ at point $F$.
  5. 5.
    Draw circle with center $A_{i}$ and radius equal to through $A_{i+1}F$. This circle intersect line segment $A_{i}A_{i+1}$ at point $B_{i}$ such that $A_{i}B_{i}=\frac{A_{i}A_{i+1}}{2\left(1+\cos^{2}{\frac{\pi}{n}}\right)}$.
The proof follows from the following arguments. Note first that $|A_{i}C|=a\cos{\frac{\pi}{n}}$, where as before $a=|E_{i}A_{i}|$. Then $|E_{i}D|=a\cos^{2}{\frac{\pi}{n}}$ and therefore $|A_{i+1}D|=a\left(1+\cos^{2}{\frac{\pi}{n}}\right)$. Since $|A_{i+1}E_{i}|=a$,$|A_{i+1}G|=a$. Also from similarity of $\triangle A_{i+1}E_{i}F$ and $\triangle A_{i+1}DG$ we obtain that $\frac{|A_{i+1}E_{i}|}{|A_{i+1}D|}=\frac{|A_{i+1}F|}{|A_{i+1}G|}$, which simplifies to $|A_{i+1}F|=\frac{a}{1+\cos^{2}{\frac{\pi}{n}}}$.
Note also that if regular polygon $A_{0}A_{1}A_{2}\ldots A_{n-1}$ and point $B_{i}$ with specific $x=x_{0}$ $(0<x_{0}<a,\ x_{0}\neq\frac{a}{1+\cos{\frac{\pi}{n}}})$ is given, then one can construct using the same instruments another point $B^{\prime}_{i}$ with $x=x_{1}=|A_{i}B^{\prime}_{i}|$, such that $f(x_{0})=f(x_{1}).$ Indeed, $x_{1}$ is the second solution of quadratic equation $\frac{x(a-x)\sin{\frac{2\pi}{n}}}{2a-2x\sin^{2}{\frac{\pi}{n}}}=f(x_{0})$ (the first solution is $x=x_{0}$), all of the coefficients of which are constructible. For $x_{0}=\frac{a}{1+\cos{\frac{\pi}{n}}}$, one can check that $x_{1}=x_{0}$.
Figure 8: Construction of $B_{i}$ such that $A_{i}B_{i}=\frac{A_{i}A_{i+1}}{2\left(1+\cos^{2}{\frac{\pi}{n}}\right)}$.
5 Appendix B
In this part of the paper a new proof of generalized Maclaurin-Braikenridge’s theorem for $n$-gons $(n\geq 3)$ will be given. It is based on the method of mathematical induction for $n$. First, we need to prove the following lemma.
Lemma 5.1.
Let fixed points $B$ and $F$, and moving point $E$ be on a given conic $\omega$. Let $P$ and $l$ be a given point and a given line on the plane of $\omega$. Let $FE$ intersect $l$ at point $K$, and $PK$ intersect $BE$ at $T$. Then the locus of point $T$ is a conic passing through points $P$ and $B$ (see Fig. 9).
Figure 9: Lemma 5.1.
Proof.
Consider 2 cases.
Case 1. Suppose that line $l$ does not intersect conic $\omega$. Then we can apply a projective transformation mapping $l$ to infinite line and then an affine transformation mapping the image of conic $\omega$ to a circle. Let us use the same letters for the new objects. When $E$ changes on circle $\omega$, $\angle FEB$ does not change. Since $K$ is now an infinite point on infinite line $l$, $PK\parallel FK$, and therefore, $\angle PTB=\angle FEB=\textnormal{const}$. So, locus of $T$ is a circle. This means that inverse of affine and projective transformations of the locus of $T$ will be a conic.
Figure 10: Proof of Lemma 5.1.
Case 2. Suppose now that $l$ intersects $\omega$ at points $A$ and $D$ as in Fig. 10. Denote $C=PF\cap\omega$, $L=PF\cap BE$, $M=AC\cap BD$, $Q=AE\cap FD$, and $N=AT\cap PD$. By Pascal’s theorem, since hexagon $ACFDBE$ is inscribed into conic $\omega$, points $M,L,Q$ are collinear. By Desarques’ theorem, since $\triangle AET$ and $\triangle DFP$ are perspective with respect to point $K$ ($AD$, $EF$, and $TP$ are concurrent), these triangles are also perspective with respect to a line. Therefore points $M$, $L$, and $N$ are collinear. Then by Pascal’s theorem, hexagon $ACPDBT$ is inscribed into a conic. Since points $A,\ C,\ P,\ D,\ B$ are fixed, the locus of point $T$ is a conic. ∎
We are ready to state and prove the following general result.
Theorem 5.2.
Let $l_{0},l_{1},\ldots,l_{n-1}$ be $n$ lines, and $C_{0},C_{1},\ldots,C_{n-1}$ be $n$ points in general position $(n\geq 3)$. Let changing points $B_{i}\in l_{i}$ $(i=0,1,\ldots,n-1)$ and $B_{n}\in l_{0}$ be chosen so that $C_{i}\in B_{i}B_{i+1}$ $(i=0,1,\ldots,n-1)$. Then the locus of point $X_{n}=B_{0}B_{1}\cap B_{n}B_{n-1}$ is a conic passing through $C_{0}$ and $C_{n-1}$.
Proof.
We saw in Section 3 that the case $n=3$ is the converse of Pascal’s theorem and therefore it is easily proved [10], p. 85. Suppose that the claim is true for $n=k$ $(k\geq 3)$, which means that the locus of $X_{k}=B_{0}B_{1}\cap B_{k}B_{k-1}$ is a conic $\omega_{k}$ passing through $C_{0}$ and $C_{k-1}$. If we take in Lemma 5.1 $E=X_{k-1}$, $F=C_{0}$, $B=C_{k-1}$, $K=B_{k}$, $\omega=\omega_{k}$, $P=C_{k}$, $T=X_{k}$, $l=l_{k}$, then we obtain that the locus of $X_{k}$ is a conic passing through $C_{0}$ and $C_{k}$. ∎